OpenStudy (anonymous):

5x^(1/2)-9x^(1/4)+3=0

OpenStudy (unklerhaukus):

solve for x?

OpenStudy (anonymous):

find the real solutions

OpenStudy (aravindg):

heres what wolf has to say http://www.wolframalpha.com/input/?i=5x%5E%281%2F2%29-9x%5E%281%2F4%29%2B3%3D0

OpenStudy (anonymous):

First Solve \[5 u^2-9 u+3=0 \] then \[ x =u^4 \]

OpenStudy (anonymous):

\[ x= \frac{9 \left(239-51 \sqrt{21}\right)}{1250} \\ x=\frac{9 \left(239+51 \sqrt{21}\right)}{1250} \\ \]